4. Classify each of the following reactions as one of these four types:

• spontaneous at all temperatures

• not spontaneous at any temperature

• spontaneous below a certain temperature but not above

• spontaneous above a certain temperature but not below

(a) CO (g) + 3 H2 (g) → CH4 (g) + H2O (g); ∆H = -206.1 kJ; ∆S = -214.6 J/K

(b) AgClO3 (s) + CH4 (g) → AgCl (s) + 2 H2O (g) + CO (g); ∆H = -616 kJ; ∆S = 343 J/K

(c) CaS (s) + H2S (g) → CaH2 (s) + 2 S (s); ∆H = 317 kJ; ∆S = -157 J/K

(d) 2 NH3 (g) → N2 (g) + 3 H2 (g); ∆H = 92.2 kJ; ∆S = 198.8 J/K
Answer #1

Spontaneity of a reaction can be evaluated by calculating the change in free energy of the reaction using the relation:

\DeltaG = \Delta H - T\DeltaS

for a spontaneous reaction \Delta G <0

thus temperature for the reaction to be spontaneous, \Delta H - T\DeltaS <0

or T>\DeltaH /\DeltaS

a) CO (g) + 3 H2 (g) → CH4 (g) + H2O (g); ∆H = -206.1 kJ; ∆S = -214.6 J/K

Given that \Delta H = -206.1 kJ and \Delta S = -214.6 J/K = -0.2146 kJ/K

\DeltaG = -206.1 - T*(-0.2146)

\DeltaG = -206.1 + T*0.2146

thus \Delta G will be < 0 for 206.1>T*0.2146

or T< 960 K

Thus the reaction will be spontaneous for temperature below 960 K

(b) AgClO3 (s) + CH4 (g) → AgCl (s) + 2 H2O (g) + CO (g); ∆H = -616 kJ; ∆S = 343 J/K

Given that \Delta H = -616 kJ and \Delta S = 343 J/K = 0.343 kJ/K

\DeltaG = -616 - T*(0.343)

\DeltaG = -206.1 - T*0.343

since T> 0

thus \Delta G will be < 0 for all values of T

Therefore reaction is spontaneous at all temperatures

(c) CaS (s) + H2S (g) → CaH2 (s) + 2 S (s); ∆H = 317 kJ; ∆S = -157 J/K

Given that \Delta H =317 kJ and \Delta S = -157 J/K = -0.157 kJ/K

\DeltaG = 317 - T*(-0.157)

\DeltaG = 317 + T*0.157

as T > 0. thus \Delta G >0 at all temperatures

Thus the reaction will be non-spontaneous at all temperatures.

(d) 2 NH3 (g) → N2 (g) + 3 H2 (g); ∆H = 92.2 kJ; ∆S = 198.8 J/K

Given that \Delta H =92.2 kJ and \Delta S = 198.8 J/K = 0.1988 kJ/K

\DeltaG = 92.2 - T*(0.1988)

\DeltaG = 92.2 - T*0.1988

thus \Delta G will be < 0 for 92.2<T*0.157

or T> 587 K

Thus the reaction will be spontaneous for temperature above 587 K